OpenStudy (jonathan34):

Evaluate the expression 43 ÷ (7 − 3) × 2.

OpenStudy (jonathan34):

will fan and medal

OpenStudy (jonathan34):

thx

OpenStudy (mathmale):

jon: You need to know how to apply the PEMDAS rules yourself. Do you know them? If so, what's the first step you'd take in solving this problem? If not, better look up PEMDAS now, and take notes on it for future reference.

OpenStudy (kky10997):

\[\frac{ 43 }{ ( 7 -3) }\times2=\frac{ 43 }{ 4 }\times2=\frac{ 43 }{ 2}\]